Solve the valid solutions of the following equations providing your values for some of them. Write your answers in solution set form. Identify the extraneous solution, if there is.


x 1 8

------- - -------- = ---------

x + 2 x - 2 x^2 - 4


1
Expert's answer
2021-05-17T17:09:17-0400
"\dfrac{x}{x+2}-\dfrac{1}{x-2}=\dfrac{8}{x^2-4}"

"x(x-2)-(x+2)=8, x \not=\pm2"

"x^2-2x-x-2-8=0"

"x^2-3x-10=0"

"(x+2)(x-5)=0"

"x_1=-2, x_2=5"

But "x \not=\pm2"

The solution "x=-2" is extraneous and is not valid.


"\{5\}"